# {dedication}

### v1

@@ 1 Work Completed to Date
    Insights into the strong scaling characteristics of GNN training.
SMASH: Scalable SPGEMM Kernel Implementation:
The thesis also presents Sparse Matrix Atomic Scratchpad Hashing (SMASH), a scalable Sparse Matrix-Matrix Multiply (SpGEMM)kernel implementation designed for a custom graph accelerator. This part of the research focuses on:
@@
    Developing three iterative optimizations for SPGEMM, leveraging the unique architectural features of the graph accelerator.
@@
    Implementing atomic hashing optimization, tokenization, and pipelining of SMASH kernels.
@@
    Achieving considerable speedups, with averages of 1.20×1.20$\times$1.20 ×, 1.29×1.29$\times$1.29 ×, and 1.04×1.04$\times$1.04 × over MKL single socket, MKL dual-socket, and GPU A100 hardware, respectively.
2 Plans to Complete The Thesis
A systematic plan is laid out to complete the thesis, including motivation and details for the remaining topics that will result in the efficient acceleration of both Sparse Matrix-Matrix Multiplication (SpGEMM) and dense matrix computation that are inherent in graph neural networks (GNNs). The thesis completion plan is as follows:
  1. 1.
    Identifying Bottlenecks of Dense Workloads: The first phase involves a thorough investigation into the performance bottlenecks associated with dense workloads in graph computing. This task is crucial for understanding the limitations of current architectures and algorithms in handling dense matrix computations, which are integral to GNN processing. The insights gained from this analysis will directly inform the design and optimization strategies for the proposed GNN accelerator.
  2. 2.
    Cycle-Accurate Simulator Implementation: Subsequently, the development of a cycle-accurate simulator is planned. This simulator will play a pivotal role in modeling and evaluating the performance of the proposed graph accelerator architecture under various computational loads. By accurately simulating the behavior of the hardware, this tool will enable precise tuning of the architectural features and provide a platform for accelerator design for GNN workloads.
  3. 3.
    SpGEMM and Dense GEMM Accelerator Design: Following the simulator implementation, the focus will shift to the design of the SpGEMM and Dense GEMM accelerator. This step is central to the thesis, as it involves creating a specialized hardware solution capable of efficiently executing both sparse and dense matrix operations. The design will leverage the findings from the initial bottleneck analysis and the simulator tests to ensure that the accelerator is optimized for the specific demands of GNN computations.
@@ 2 Plans to Complete The Thesis
    Evaluation of GNN Accelerator: After the accelerator design is completed, it will undergo a comprehensive evaluation phase. This evaluation will assess the accelerator’s performance, scalability, and efficiency in executing GNN workloads. The aim is to validate the effectiveness of the designed accelerator in accelerating graph computing tasks and to benchmark its performance against existing solutions.
The timeline for the final thesis is as follows:
Task Time
Identifying bottlenecks of dense workloads 1 month
Cycle-accurate simulator implementation 2 months
SpGEMM Dense GEMM accelerator design 1 month
Evaluation of GNN accelerator 1 month
Final thesis write-up 1 month
Table 1: Plan for Completing the Thesis

### v2

@@ 4 Contributions of this Dissertation
    An in-depth evaluation of SpGEMM optimizations on a custom accelerator, showcasing the potential of hardware-specific adaptations to elevate GNN processing speeds.
  5. 5.
    The proposal of a dynamically reseeding hash-based mapping algorithm tailored for GNN workloads, which optimizes computation distribution and efficiency in hardware accelerators designed for GNNs.
  6. 6.
    The creation of NeuraSim, a cycle-accurate simulator for CGRA architectures, facilitating precise analysis and optimization of GNN accelerator designs.
  7. 7.
    A comprehensive chip power and area analysis of the NeuraChip GNN accelerator, providing valuable metrics for assessing the viability and efficiency of GNN-specific hardware solutions.
Together, these contributions not only provide a path towards optimized GNN processing but also equip the research community with the tools and methodologies necessary for continued innovation in the acceleration of graph neural networks.
5 Future Work
The primary objective of this dissertation was to advance the field of Graph Neural Network acceleration through a comprehensive exploration of benchmark suites, algorithmic optimizations, and hardware solutions. Despite the significant strides made in this research, the domain of GNN acceleration remains vast, with ample opportunities for further exploration and innovation. The following outlines potential directions for future research, building upon the foundational work presented in this thesis:
@@ 5 Future Work
    Enhanced Integration of GNN Workloads with Emerging Hardware Accelerators: The current landscape of hardware accelerators, including but not limited to NVIDIA’s Tensor Cores [123] and Google’s TPUs [99, 107], presents a promising avenue for improving GNN performance and energy efficiency. Future work could explore deeper integrations with these accelerators, adapting GNN algorithms to leverage their specific architectural advantages more effectively.
@@
    Advanced Memory Technologies for GNN Acceleration: The evolution of memory technologies, such as in-memory processing [85] and near-memory processing [95], offers new possibilities for enhancing GNN execution. Investigating holistic designs that incorporate these advanced memory solutions could lead to significant improvements in GNN processing speed, energy efficiency, and overall system performance.
@@
    Scalability and Efficiency in Multi-GPU Systems: While this dissertation addressed multi-GPU benchmarking and architectural implications for GNNs, scaling performance linearly with the addition of GPUs remains a challenge. Future research could focus on developing new techniques for GPU-level cooperative thread array (CTA) [93] scheduling and thread migration[42] to minimize performance overhead in multi-GPU systems.
@@
    Dynamic Scheduling and Control Flow Optimization for GNN Workloads: Addressing the limitations in control flow transition between CPUs and GPUs is crucial for optimizing the performance of GNN workloads [98]. Investigating hardware-based GPU schedulers integrated into CPU cores could reduce kernel-launch overhead and memory synchronization challenges, enabling more efficient CPU-GPU collaboration.
  5. 5.
    Computing Capabilities in Network Devices for Distributed GNN Processing: As network switches evolve to incorporate computing capabilities, there exists an opportunity to offload certain GNN processing tasks to these devices, potentially reducing data movement costs and improving the efficiency of distributed GNN training and inference [180].
  6. 6.
    Exploration of Novel Computing Devices for GNN Acceleration: The advent of novel computing devices extends the horizon for GNN acceleration. Future research should consider incorporating a broader array of devices, including dedicated accelerators, non-traditional memory devices, and smart network components, to achieve comprehensive performance, energy, reliability, and security enhancements.
Building on the contributions of this dissertation, these future research directions aim to further push the boundaries of GNN acceleration, addressing the complex and evolving challenges inherent in graph-based computations and fostering the development of more efficient, scalable, and adaptable GNN processing systems.
Biography
Kaustubh Shivdikar
Kaustubh Shivdikar was born in Mumbai, India, on December 5, 1994. He obtained his Bachelor of Science degree in Electrical Engineering from the Veermata Jijabai Technological Institute, University of Mumbai, in 2016. He went on to receive his Master of Science and Doctor of Philosophy degrees in Electrical and Computer Engineering from Northeastern University, Boston, USA, in May 2020 and May 2024, respectively. His Ph.D. research was supervised by Dr. David Kaeli at the Northeastern University Computer Architecture Research (NUCAR) Laboratory. Kaustubh is a member of IEEE and ACM. His research fields encompass Computer Architecture Simulator Design, Graph Neural Network Accelerators, Sparse Matrix Accelerators, and Homomorphic Encryption Accelerators. [^1] https://wiki.kaustubh.us
Index
Enabling Accelerators for Graph Computing
Kaustubh Shivdikar